### v1

# Volumetric Semantically Consistent 3D Panoptic Mapping

Abstract
We introduce an online 2D-to-3D semantic instance mapping algorithm aimed at generating comprehensive, accurate, and efficient semantic 3D maps suitable for autonomous agents in unstructured environments. The proposed approach is based on a Voxel-TSDF representation used in recent algorithms. It introduces novel ways of integrating semantic prediction confidence during mapping, producing semantic and instance-consistent 3D regions. Further improvements are achieved by graph optimization-based semantic labeling and instance refinement. The proposed method achieves accuracy superior to the state of the art on public large-scale datasets, improving on a number of widely used metrics. We also highlight a downfall in the evaluation of recent studies: using the ground truth trajectory as input instead of a SLAM-estimated one substantially affects the accuracy, creating a large gap between the reported results and the actual performance on real-world data. The code is available at: https://github.com/y9miao/ConsistentPanopticSLAM.
I INTRODUCTION
Autonomous agents operating in unstructured environments require a comprehensive understanding of their surroundings to make informed decisions. In addition to perceiving the geometry of the environment, this includes understanding semantic categories and recognizing individual instances of objects within the scene. Robust perception of the surrounding complex space enables the agent to discover, segment, track, and reconstruct objects at the instance level to guide their subsequent actions. However, real-world scenarios exhibit considerable variability in object appearance, shape, and location, posing a challenge to robotic perception.
3D semantic-instance mapping is the problem of simultaneously detecting and segmenting objects with their semantic and instance labels to generate a 3D semantic map of the surrounding environment. There are two popular approaches to this problem: 3D-to-3D, which analyzes input 3D point clouds – in this scenario the environment is previously reconstructed; and 2D-to-3D, which analyzes a set of 2D images, further projecting and aggregating predictions on the 3D map – here reconstruction and semantic mapping is done synchronously. In practice, 3D-to-3D approaches often suffer from a lack of large-scale training data and speed issues, i.e., 3D data is more complex than 2D, which makes processing and analyzing 3D point clouds computationally intensive. 2D-to-3D algorithms leverage state-of-the-art 2D semantic-instance segmentation algorithms hence resolving some of the issues with 3D-to-3D approaches. However, having only 2D observations of the 3D instances, such methods often produce a noisy segmentation of the environment due to the limited perspective of 2D images and the noise in semantic detections per frame and camera calibration.
In this paper, we address the problem of noisy semantic detections by implementing semantic consistency and further regularization on semantics and instances. Specifically, we propose a 2D-to-3D semantic and instance mapping algorithm that achieves accuracy superior to state-of-the-art 2D-to-3D semantic instance segmentation methods on public large-scale datasets. Our approach is inspired by Voxblox++ [1], which we improve upon with the following contributions:
  1. 1.
    A new way of employing 2D semantic prediction confidence in the mapping process (Section III-A).
  2. 2.
    Developing a novel method to segment semantic-instance consistent surface regions, i.e., super-points (Section III-B).
  3. 3.
    A new graph optimization-based semantic labeling and instance refinement algorithm (Sections III-C & III-D).
II RELATED WORK
The task of 3D semantic instance mapping has a number of real-world applications in computer vision and robotics, allowing autonomous agents to understand their surroundings. Methods used for this task can be divided into two main categories based on their input.
3D-to-3D approaches primarily analyze input 3D point clouds to discern instances directly in 3D. Bottom-up approaches [2, 3, 4] form object instances by grouping points based on their local affinities. In contrast, top-down approaches [5, 6] initiate the process by identifying object bounding boxes and subsequently segmenting them into foreground and background regions. Another successful methodology is the voting-based approaches [7, 8, 9, 10, 11], where points nominate or ”vote” for object centers, followed by a clustering process of the center votes. Recent Transformer-based strategies [12, 13] represent objects using learned instance queries, correlating them with per-point features to create instance masks.
While these approaches can provide highly accurate results since they have access to the entire geometry of the scene and the objects, they often require a significant amount of computational resources and are challenging to scale to large scenes. In addition, they require a 3D map to be previously reconstructed, hence resulting in asynchronous methods.
2D-to-3D approaches take 2D images as input and progressively reconstruct the scene in a semantic and instance-aware manner. This approach has received increasing attention, especially in the robotics field, due to its capability to be real-time and incremental – factors that are indispensable for applications such as navigation and path planning.
Among the representative techniques are Voxblox++ [1] and PanopticFusion [14]. Voxblox++ [1] employs both semantic instance and geometric segmentation, converting 2D frames into instance-labeled 3D segments. The process involves the incremental registration of geometric segments detected in depth frames onto global 3D counterparts, subsequently integrating semantic and instance labels within these global segments. Our method is inspired by [1], but it offers substantial enhancements, particularly in terms of semantic accuracy. In [14], panoptic segmentation [15] is applied to extract semantic instances for “things” and semantic labels for collective objects (“stuff”), which help improve the quality of semantic-instance segmentation. SceneGraphFusion [16] builds a scene graph with geometric segments and then applies a graph neural network to predict semantic and instance labels. However, SceneGraphFusion utilizes depth images synthetically rendered directly from a reconstructed 3D mesh, hence using ground truth data. The recent INS-CONV [17] combines the advantages of 2D-3D and 3D-3D procedures. While it builds the 3D model of the environment incrementally from RGB-D frames, it performs segmentation directly in the 3D space. This, however, comes at the cost of requiring 3D ground truth annotations for training.
Figure 1: In III-A, the proposed method gets an RGB-D sequence and runs panoptic and geometric segmentation that are then fused to extract 3D surfaces with panoptic labels. In III-B, it incrementally lifts frame-wise surfaces into a global coordinate system to create semantically consistent superpoints. In III-C, superpoint graph is constructed and updated. Semantic segmentation is performed using the graph. In III-D, the initial instance segmentation is refined by graph optimization.
Noisy 2D predictions stand as a recurrent source of errors in 2D-to-3D techniques. Both geometric [18] and panoptic segmentations [19, 15] occasionally yield over- or under-segmented scenes, leading directly to inaccuracies in the 3D segmentation. To address geometric over-segmentation, Voxblox++ [1] adopts a strategy of iteratively merging proximate 3D segments. However, this method falls short in considering semantic consistency, resulting in incorrect mergers of disparate segments. To rectify this issue, both [14] and [20] employ a computationally expensive voxel-level map regularization. Han et al. [21] make strides in the direction of leveraging spatial context, opting for regularization at the segment level. This, however, can only deal with 3D instance over-segmentation. Distinctively, our approach incorporates super-point-level regularization and refinement. This not only ensures enhanced semantic-instance segmentation but also retains a small computational footprint.
III SEMANTIC-INSTANCE MAPPING
The proposed incremental 2D-to-3D instance and semantic mapping approach, shown in Fig. 1, takes a sequence of RGB-D frames and poses estimated by SLAM pipeline, such as [22], as input, and goes through a six-step procedure for each RGB-D frame: (i) 2D panoptic-geometric segmentation (Sec. III-A); (ii) 3D superpoint segmentation (III-B); (iii) superpoint graph construction and update (III-C). The next steps run only when corresponding information is queried: (iv) semantic regularization (III-C); (v) instance refinement (III-D); and (vi) mesh generation as in [1].
III-A 2D Panoptic-Geometric Segmentation
Panoptic-geometric segmentation (Fig. 1 III-A), is applied on each RGB-D frame $F$ based on the assumption [18]:
Assumption 1
The surface of each real-world object consists of single or multiple neighboring convex surfaces which share the same semantic-instance label.
First, panoptic segmentation [23] on the RGB image provides a set of semantic instance masks and geometric segmentation [18] on the depth image extracts a set of convex segments. Intersections between panoptic and geometric segmentation masks are used to generate a set of surface masks $\mathcal{F}_{s}=\{s_{i}\}$ per frame $F$. In this way, each surface mask $s_{i}$ is frame-wise uniquely associated with a panoptic semantic instance $o_{i}$ and a geometric segment, thus being less prone to geometric and semantic under-segmentation.
Each surface $s_{i}$ is assigned with instance label $o_{i}$, semantic category $C(o_{i})$, and panoptic confidence defined as:
$P_{o_{i}}=\left\{\begin{array}[]{rcl}\text{Score}(o_{i})\in(0,1)&&{C(o_{i})\in C_{\text{Th}}},\\ 0.5&&{C(o_{i})\in C_{\text{St}}},\\ \end{array}\right.$ (1)
where $C_{\text{Th}}$ and $C_{\text{St}}$ are the sets of “thing” and “stuff” categories [23]. $\text{Score}(o_{i})$ is the confidence score for instance $o_{i}$ from [23]. Given the camera pose and depth image, the point cloud $\text{Pcl}(s_{i})$ of $s_{i}$ is projected to a unified Voxel-TSDF structure, that is incrementally built.
III-B Semantically Consistent Superpoint Segmentation
In addition to the above, each surface $s_{i}\in\mathcal{F}_{s}$ is also associated with a superpoint label $L_{S}(s_{i})$. Next, we explain what is a superpoint and how to compute the association between a surface and a superpoint.
Based on Assumption 1, we define a “superpoint” $S$ as the set of neighboring convex surfaces stemming from the same 3D object. Each superpoint has a unique identifier, or else label, $L_{S}$. This leads to the following assumption:
Assumption 2
Each real-world object consists of single or multiple spatially connected superpoints with the same semantic-instance labels.
Based on Assumptions 1 and 2, we interpret surface $s_{i}\in\mathcal{F}_{s}$ as a partial observation of a superpoint $S$ in the global 3D space. Each surface $s_{i}$ is assigned to a superpoint by calculating the overlap $\Pi_{s}$ between $s_{i}$ and candidate superpoints in the 3D space (i.e., the number of overlapping voxels in the Voxel-TSDF structure). Specifically, for each point $p_{i}$ of $\text{Pcl}(s_{i})$, we query the superpoint label $L_{S}$ in voxel $V(p_{i})$ and then update the overlap count incrementally as $\Pi_{s}(L_{S}(V(p_{i})),s_{i}):=\Pi_{s}(L_{S}(V(p_{i})),s_{i})+1$. Surface $s_{i}$ is assigned to superpoint $L_{S}$ if the overlap is large enough, controlled by a hyper-parameter. Otherwise, a new label $L_{S_{\text{new}}}$ is created for $s_{i}$.
To ensure that assumption 2 is met, we maintain $\Pi_{V}(L_{S},V_{i})$, the count of occurrences of superpoints $L_{S}$ in a voxel $V_{i}$. The final superpoint label $L_{S}^{*}$ for $V_{i}$ is
$L_{S}^{*}(V_{i})=\arg\max_{L_{S}}\Pi_{V}(L_{S},V_{i}).$ (2)
i.e., a voxel is assigned to the superpoint with the most votes. We use the above equation when incorporating a new surface $s_{i}$ in the global map and querying the current superpoint label $L_{S}$ in voxel $V(p_{i})$.
To address the over-segmentation of superpoints, a pairwise (voxel) overlap score $M(L_{S_{1}},L_{S_{2}})$ is maintained and updated for each surface $s_{i}$ as follows:
$\small M(L_{S_{1}},L_{S_{2}})=\sum_{t=0}^{n}\sum_{s_{i}\in\mathcal{F}_{s}}\llbracket S_{1}\in L_{N}(s_{i})\wedge S_{2}\in L_{N}(s_{i})\rrbracket,$ (3)
[…]
  1. 1.
    Spatially connected: $M(L_{S_{1}},L_{S_{2}})>\theta_{\text{merge}}$.
  2. 2.
    Semantically consistent: $C_{S}(L_{S_{1}})$ equals to $C_{S}(L_{S_{2}})$ or at least one of $C_{S}(L_{S_{1}}),C_{S}(L_{S_{2}})$ equals to $C_{0}$, where $C_{S}$ is the semantic label of a superpoint, determined by Eq. 4 and $C_{0}$ is the background semantic label.
Figure 2: Superpoint segmentation without (left) and with (right) semantic consistency as proposed in Section III-B
Considering semantic consistency helps ensure that Assumption 2 holds and, thus, achieves higher superpoint segmentation accuracy than [1, 21]. An example is shown in Fig. 2.
III-C Semantic Segmentation via Super-Point Graph
To address that 2D instance labels are often inconsistent across frames since inference happens frame-wise, prior work [1, 21] apply a data association step to find correspondences of 2D instances and 3D instance labels. Essentially, this process counts observations of semantic-instance labels and assigns superpoints to the labels with the maximum count. However, this method has its pitfalls. It simplistically assigns an instance to the superpoint that has the greatest overlap. As a result, it tends to miss fine details, leading to under-segmentation. For instance, smaller objects near larger ones might be misidentified as a part of the bigger entity (like mistaking a mug as a part of a table). Conversely, it is prone to over-segmentation, especially when there is minimal overlap between the frame-wise segments and superpoints, often due to inaccuracies in pose or depth measurements. This often leads to superfluous instance labels in practice.
To tackle the aforementioned challenges, we leverage a superpoint graph to aggregate frame-wise semantic-instance-spatial data, which aids in determining the semantic-instance label assigned to a superpoint. Let us construct a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with superpoints as vertices ($L_{S}\in\mathcal{V}$). We connect two vertices by edge $\epsilon^{C}=(L_{S_{1}},L_{S_{2}})\in\mathcal{E}$ if the spatial-panoptic confidence $P_{\epsilon}^{C}(L_{S_{1}},L_{S_{2}})>0$. Confidence $P_{\epsilon}^{C}(L_{S_{1}},L_{S_{2}})$ denotes how likely two superpoints $L_{S_{1}},L_{S_{2}}$ share the same instance label with semantic class $C$ and are spatially connected. Each node has its semantic confidence $P_{\nu}^{C}(L_{S})$ representing how likely $L_{S}$ belongs to semantic $C$.
Confidence calculation. Here, we describe how confidence $P_{\nu}^{C}$ and $P_{\epsilon}^{C}$ is calculated. For frame $F$, we ray-cast the pixels of each instance $o_{i}$ into the 3D superpoint space. The confidence being assigned to class $C(o_{i})$ in superpoint $S$ comes from aggregating the voxels of $S$ that intersect the rays. Precisely, $P_{\nu}^{C}(L_{S})=P_{\nu}^{C}(L_{S})+P_{o_{i}}\cdot N^{S}_{o_{i}}$, $N^{S}_{o_{i}}=\sum_{V\in\text{voxels}_{L_{S}}}\llbracket\text{intersect}(V,\text{rays}(o_{i}))\rrbracket$. If two superpoints are hit by rays of $o_{i}$, pairwise confidence is updated: $P_{\epsilon}^{C(o_{i})}(L_{S_{1}},L_{S_{2}})=P_{\epsilon}^{C(o_{i})}(L_{S_{1}},L_{S_{2}})+P_{o_{i}}\cdot P_{spatial}\cdot min(N^{S_{1}}_{o_{i}},N^{S_{2}}_{o_{i}})$, $P_{spatial}$ is the spatial confidence determined by spatial distance between $L_{S_{1}}$ and $L_{S_{2}}$. The initial guess of semantic labels to superpoints is assigned and updated during incremental mapping:
$C(L_{S})=\max_{C}P_{\nu}^{C}(L_{S}),$ (4)
Graph optimization. However, such assignment is based only on individual observations of each superpoint. To further optimize this assignment, we propose a graph-based approach employing the $\alpha-\beta$ swap algorithm [24].
Given graph $\mathcal{G}$ constructed as described previously, we apply the $\alpha-\beta$ swap algorithm, minimizing energy
$\small E(C)=\sum_{L_{S}\in\mathcal{V}}\psi_{u}(C(L_{S}))+\sum_{(L_{S_{1}},L_{S_{2}})\in\mathcal{E}}\psi_{p}(C(L_{S_{1}}),C(L_{S_{2}})),$ (5)
where unary potential $\psi_{u}(C(L_{S}))$ is defined with negative logarithm of a probability distribution $p_{C(L_{S})}(C)$ over semantic labels as follows:
$\begin{split}&p_{C}(C(L_{S}))=\frac{P_{\nu}^{C(L_{S})}(L_{S})}{\sum_{C}P_{\nu}^{C}(L_{S})},\\ &\psi_{u}(C(L_{S}))=-\log p_{C}(C(L_{S})).\end{split}$ (6)
Binary potential $\psi_{p}(C(L_{S_{1}}),C(L_{S_{2}}))$ is as follows:
$\small\begin{split}&k(L_{S_{1}},L_{S_{2}},C(L_{S_{1}}),C(L_{S_{2}}))=\frac{{P_{\nu}}^{C(L_{S_{1}})}({L_{S_{1}}}){P_{\nu}}^{C(L_{S_{2}})}({L_{S_{2}}})}{\sum_{C(L_{S})}P_{\epsilon}^{C(L_{S})}(L_{S_{1}},L_{S_{2}})},\\ &\psi_{p}(C(L_{S_{1}}),C(L_{S_{2}}))=\mathbb{1}_{C(L_{S_{1}})\neq C(L_{S_{2}})}K_{C}\exp{\left(-\frac{k(*)}{2\theta^{2}}\right)},\end{split}$ (7)
where $k(*)$ is $k(L_{S_{1}},L_{S_{2}},C(L_{S_{1}}),C(L_{S_{2}}))$ and it indicates how likely $L_{S_{1}}$ and $L_{S_{2}}$ stem from different semantic-instance labels. It is straightforward to see that the energy is submodular and semi-metric and, thus, can be efficiently minimized by the $\alpha-\beta$ swap algorithm. Note that to speed up processing, semantic regularization runs only in case of semantic query, i.e. mesh generation or query for planning.
TABLE I: The mAP50 scores (higher is better) on 10 scenes of the SceneNN dataset [25] using the GT (top) and SLAM-estimated [22] (bottom) camera poses. Results on GT are from [20]. The SLAM part shows methods with public code.
Method / Sequences 011 016 030 061 078 086 096 206 223 255 Average
GT trajectory Voxblox++ [1] 175.0 48.2 62.4 66.7 155.8 20.0 34.6 79.6 43.8 175.0 56.1
Han et al. [21] 165.8 50.0 66.6 43.3 100.0 56.9 22.8 92.1 46.7 133.0 57.7
Wang et al. [26] 162.2 43.0 60.7 36.3 149.3 45.8 32.7 46.0 46.6 156.4 47.9
Li et al. [27] 178.6 25.0 58.6 46.6 169.8 47.2 26.7 78.0 45.8 175.0 55.1
Mascaro et al. [20] 100.0 75.0 72.5 50.0 150.0 50.0 51.3 74.1 45.8 100.0 66.9
INS-CONV [17] 100.0 62.0 83.4 69.8 193.7 60.0 57.6 56.7 78.6 100.0 76.2
Ours 100.0 73.3 91.7 62.4 187.5 61.7 66.7 83.3 60.0 100.0 78.7
SLAM Voxblox++ [1] 161.5 38.9 50.0 58.4 144.3 16.4 27.6 48.7 40.7 133.6 42.0
Han et al [21] 153.4 43.2 50.0 37.6 175.6 48.2 13.4 57.8 44.1 124.7 44.8
INS-CONV [17] 175.0 46.7 56.4 57.1 183.7 22.4 48.1 28.1 65.2 150.0 53.3
Ours 175.0 56.7 72.3 62.4 168.9 55.6 33.2 40.8 60.0 163.3 58.8
III-D Instance Refinement
This section focuses on improving instance segmentation. Existing methods like [1, 21] commonly face issues of under-segmentation, where two or more instances with the same semantic label get merged. This merging is often caused by under-segmentation in 2D frames or 3D superpoints. Conversely, over-segmentation can occur because of unreliable data association from noisy camera poses and depth measurements. To tackle these challenges, we employ instance refinement on the instance labels. We run refinement separately for each semantic class as the semantic labels were determined previously. This entails running the refinement on the superpoint set associated with a particular semantic label.
Suppose that we are given a semantic label $C$ and the set $\mathcal{V}_{C}$ of superpoints assigned to $C$, and initial instance guess for superpoints $\mathcal{O}_{C}=\{O_{L_{S}}\}$, where ${L_{S}}\in\mathcal{V}_{C}$. The goal is to optimize $\mathcal{O}_{C}$. For each unique instance label $O^{C}_{i}\in\mathcal{O}_{C}$, we define the instance confidence $P_{O}(O^{C}_{i})$, using the graph from the previous section, as the largest edge confidence given the semantic class as follows:
$P_{O}(O^{C}_{i})=\max_{(L_{S_{1}},L_{S_{2}})\in\mathcal{E}_{C}(O^{C}_{i})}P_{\epsilon}^{C}(L_{S_{1}},L_{S_{2}}),$ (8)
where $\mathcal{E}_{C}(O^{C}_{i})$ is the set of edges representing semantic class $C$ of the superpoints in $O^{C}_{i}$.
We iterate over nodes $L_{S}\in\mathcal{V}_{C}$, and detach node $L_{S}$ from its initial guess $O_{L_{S}}$ if it has low edge confidence within $O_{L_{S}}$: $P^{C}_{L_{S}}(O_{L_{S}})<\theta^{C}_{D}P_{O}(O_{L_{S}})$, where $P^{C}_{L}(O)$ is the largest edge confidence ${P_{\epsilon}}^{C}(L,L_{S})$ between superpoint $L$ and other superpoints $L_{S}$ stemming from instance $O$. We select the node with the highest confidence $P_{L}^{C}(L_{S})$ among the detached nodes and, progressively, the less confident ones. For each detached node $L_{S}$, we iterate through its neighbors and assign $O_{\text{neigh}}$ to it, if $P^{C}_{L_{S}}(O_{\text{neigh}})>\theta^{C}_{O}P_{O}(O_{\text{neigh}})$ and $P^{C}_{L_{S}}(O_{\text{neigh}})>\theta^{C}_{L}P_{L}^{C}(L_{S})$. If no such neighbors exist, $L_{S}$ will be assigned to a new instance label $O_{\text{new}}$. Neighbors of node $L_{S}$ are defined as nodes connected with $L_{S}$ in a spatial-panoptic manner: $P_{\epsilon}^{C}(L_{S},L_{\text{neigh}})>0$.
IV EXPERIMENTS
We evaluate on the SceneNN [25] and ScanNet v2 datasets [28], containing RGB-D sequences, ground truth (GT) 3D meshes, camera poses, and semantic instance annotations. We compare the proposed method with multiple state-of-the-art frameworks [20, 21, 1, 26, 27, 17]. We use four metrics to compare accuracy, computed by thresholding the intersection over union (IoU) at thresholds $0.5$ and $0.75$: (i) mean average precision (mAP), (ii) total number of true positive predicted instances ($N_{\text{TP}}$), (iii) panoptic quality (PQ) [15], and the (iv) Intersection over Union between GT instance and segmented superpoints ($\text{IoU}_{L_{S}}$), defined as $\text{IoU}_{L_{S}}=\frac{1}{\left|O\right|}\sum_{o\in O}\sum_{L_{S}}\text{IoU}(L_{S},o)$, where $O$ represents the set of all GT instances. Metric $\text{IoU}_{L_{S}}$ measures the segmentation accuracy of superpoints, and is only for methods performing superpoint segmentation and having public source code [1, 21]. We use the [23] panoptic segmentation network for all methods.
As per standard practice, we run methods using GT camera poses. Additionally, we run all approaches on poses estimated by ORB-SLAM3 [22] to demonstrate the significant performance gap between the practical setting (poses from SLAM) and the practice followed in recent papers (GT pose).
For SLAM-based experiments, we limit our evaluations to [1, 21, 17], as these are the sole methods with public code. Since [17] does not offer training code and their model has been trained on the ScanNet training set, we ensure consistency by fine-tuning or training all methods on ScanNet for the SLAM-based evaluations. Consequently, these experiments also serve as a valuable test of generalization.
[…]
(a) Comparison of methods.
Method $\text{mAP}_{50}$ $\text{mAP}_{75}$ ($\uparrow$) $N_{\text{TP}_{50}}$ $N_{\text{TP}_{75}}$ ($\uparrow$) $PQ_{50}$ $PQ_{75}$ ($\uparrow$) $\text{IoU}_{L_{S}}$ ($\uparrow$)
GT trajectory Voxblox++ [1] 33.0 24.6 253 191 21.4 15.7 64.2
Han et al. [21] 40.3 32.7 286 242 35.6 27.9 73.8
INS-CONV [17] 51.9 38.2 341 266 42.4 33.4
Ours 51.4 35.8 329 243 43.8 29.4 82.5
SLAM Voxblox++ [1] 25.7 17.6 188 169 16.2 13.0 56.8
Han et al. [21] 27.9 16.8 204 165 15.4 11.9 61.3
INS-CONV [17] 37.7 25.9 277 194 18.5 15.7
Ours 41.4 30.7 304 216 23.7 17.4 74.3
(b) Ablation study.
Method $\text{mAP}_{50}$ $\text{mAP}_{75}$ ($\uparrow$) $N_{\text{TP}_{50}}$ $N_{\text{TP}_{75}}$ ($\uparrow$) $PQ_{50}$ $PQ_{75}$ ($\uparrow$) $\text{IoU}_{L_{S}}$ ($\uparrow$)
GT trajectory 1. Ours 51.4 35.8 329 243 43.8 29.4 82.5
2. Ours w/o III-B 42.8 30.1 295 212 32.3 27.0 68.7
3. Ours w/o III-C 48.6 34.1 311 231 36.2 28.2 82.5
4. Ours w/o III-D 47.9 33.7 307 227 35.7 27.8 82.5
5. Ours w/o III-B & III-C 43.4 30.6 289 214 31.9 26.5 68.7
6. Ours w/o III-B & III-D 41.7 29.3 282 208 30.4 27.1 68.7
7. Ours w/o III-C & III-D 47.3 33.4 302 225 35.0 28.4 82.5
[…]
SLAM 1. Ours 41.4 30.7 304 216 23.7 17.4 74.3
2. Ours w/o III-B 33.0 24.2 229 182 18.5 14.8 63.7
3. Ours w/o III-C 39.1 28.9 283 209 22.6 16.9 74.3
4. Ours w/o III-D 38.5 27.5 278 201 21.7 16.2 74.3
5. Ours w/o III-B & III-C 34.9 22.2 237 176 17.6 14.6 63.7
6. Ours w/o III-B & III-D 30.7 21.5 218 172 15.9 13.7 63.7
7. Ours w/o III-C & III-D 37.4 28.3 273 205 19.8 16.1 74.3
[…]
The pipeline is implemented in C++ and runs on an Intel-12700K CPU. The RGB panoptic segmentation runs on an Nvidia RTX3090. We run [1, 21] on same machine, while [17] is run on an E5-2680 CPU and Titan Xp GPU due to library dependency issue. For all sequences in SceneNN and ScanNet v2, we use the same hyper-parameters: $\theta_{merge}=3$ (Sec. III-B), $K_{C}=15.0,\theta=0.5$ (Sec. III-C), Parameters $\theta^{C}_{O}$, $\theta^{C}_{D}$, $\theta^{C}_{L}$ (Sec. III-D) are tuned on ScanNet for 18 semantic categories. The Label-TSDF voxel size is $1$cm.
IV-A Experiments on SceneNN
As [20, 26, 27] only report mAP with IoU thresholded at 0.5 and their code is not available, we follow their evaluation setting on 10 indoor sequences from SceneNN, considering 9 object categories (bed, chair, sofa, table, books, fridge, TV, toilet, and bag), and report $\text{mAP}_{50}$. For fair comparison, the CAD post-processing from Han et al. [21] is not used since it needs a CAD object database and replaces the predicted instances with models from the database.
The results are shown in Table I. In the upper section, we evaluate methods using GT camera poses as input. The proposed method stands out, achieving the state-of-the-art average mAP50. When we switch to SLAM poses as input (lower section), there is a notable decline in the accuracy across all methods, with the average mAP50 reducing by ${\approx}20$. Even under these conditions, our algorithm remains the most accurate one. Notably, the gap between our method and the second best (INS-CONV) widens, underscoring the robustness and accuracy of the proposed approach.
[…]
TABLE III: Results on the validation set (200 scenes) of the ScanNet v2 dataset [28] using the GT (top) and SLAM-estimated [22] camera poses. The reported metrics are the class-averaged mAP, panoptic quality (PQ), and class-summed $N_{\text{TP}}$. Also, we show our results with the ground truth instance segmentation mask as an accuracy upper bound.
Method $\text{mAP}_{50}$ $\text{mAP}_{75}$ ($\uparrow$) $N_{\text{TP}_{50}}$ $N_{\text{TP}_{75}}$ ($\uparrow$) $PQ_{50}$ $PQ_{75}$ ($\uparrow$)
GT trajectory Voxblox++ [1] 47.2 23.9 2493 1408 38.6 21.7
Han et al. [21] 49.7 25.8 2586 1476 40.3 24.4
INS-CONV [17] 57.4 30.2 2917 1736 48.2 28.5
Ours 55.5 29.8 2862 1672 46.8 29.1
[…]
SLAM Voxblox++ [1] 21.9 14.1 1389 692 23.2 13.4
Han et al. [21] 25.7 15.4 1479 754 23.9 15.8
INS-CONV [17] 29.4 17.6 1764 803 25.4 16.9
Ours 32.9 20.8 1902 932 31.2 20.5
[…]
IV-B Experiment on ScanNet v2
We run experiments on the validation set of ScanNet v2 [28] consisting of 200 sequences. The results in Table III exhibit a similar trend as the ones on SceneNN. Running methods on the GT poses (upper part) leads to accurate results, with the proposed method being the second best in most metrics, trailing slightly behind INS-CONV. Using SLAM poses, as would be done in practice, significantly reduces the accuracy of all methods and changes the ranking. In this scenario, the proposed method outperforms the recent INS-CONV in all metrics, showcasing the importance of using estimated camera trajectories.
We also ran the proposed algorithm using the GT 2D instance segmentation masks to provide an upper bound on the accuracy when having perfect 2D panoptic segmentation. The accuracy significantly improves in both the GT and SLAM camera pose cases, showcasing that the proposed algorithm can be straightforwardly and substantially enhanced with the development of new panoptic networks. Let us note that we did not run other methods with GT masks, as, with this experiment, we only wanted to highlight the dependency on the panoptic segmentation network.
IV-C Ablation Study
Table II(b) displays the ablation study with components removed from the proposed method. Label ”w/o III-B” means we use the superpoint segmentation module from [21] instead of the proposed one; ”w/o III-C and w/o III-D” means no semantic regularization and no instance refinement, respectively; ”w/ GT Mask” means we use ground truth panoptic masks in III-A by rendering from GT meshes instead of using predictions from [23]. We show results both in the GT pose and SLAM cases. While the two cases have a significant performance gap, ranking the components by importance seems similar. The conclusions are as follows:
  1. 1.
[…]
  2. 2.
    Method 1 compared to 3, 1 to 4 and 1 to 7. The results show that the methods from both Sections III-B and III-D play significant roles in improving the accuracy of semantic-instance segmentation.
  3. 3.
    Method 1 compared to 8. The results shows that the 2D panoptic predictions from [23] is one of the main factors that limit our semantic segmentation accuracy. Our method improves significantly with more accurate 2D panoptic predictions as input.
IV-D Run-time and Memory Analysis
Table IV(a) shows the run-time and memory analysis of each component in the pipeline. The 2D panoptic-geometric segmentation and fusion in Section III-A takes the majority of the time and memory usage. It is important to note that this largely depends on the employed panoptic segmentation network. Thus, changing the network substantially changes the required time and memory. With limited computational resources, a lightweight panoptic segmentation (e.g., [29]) and lower voxel resolution can be used.
Table IV(b) compares the average framerates with [21, 1, 17]. Similar to [1], III-A is put into a separate thread and runs parallel to the main mapping thread to speed up. The proposed method leads to similar framerates to other methods, all running in real-time. We believe it can still be accelerated by further code optimization. Note that semantic regularization and instance refinement run only once at the end of mapping for mesh generation.
TABLE IV: Run-time and memory analysis on the 10 sequences of SceneNN dataset [25] evaluated in Table I.
(a) Execution Times of Components
Component Run-Time ($ms$) Memory Usage (MB)
Sec. III-A 216.0 (each frame) 115014 + 4359 (GPU)
Sec. III-B 170.3 (each frame) 111168.1
Sec. III-C (graph update) 127.2 (each frame) 111133.8
Sec. III-C (semantic reg.) 324.0 (only once) 111110.3
Sec. III-D (instance ref.) 119.4 (only once) 111110.1
Label-TSDF Map [1530.3, 2072.0]
(b) Average Frame-Rate Comparison
Method Voxblox++ [1] Han [21] INS-CONV [17] Ours
Framerate (Hz) 3.82 3.66 5.46 3.76
V CONCLUSIONS
In this paper, we introduce a novel 2D-to-3D semantic instance segmentation algorithm. Drawing upon a Voxel-TSDF representation, our method combines panoptic prediction confidence, semantically consistent superpoints, and graph-optimized semantic labeling and instance refinement. These components allow the proposed algorithm to outperform current state-of-the-art techniques on public and large-scale datasets. We achieve this superior accuracy while maintaining real-time processing speeds comparable to less accurate alternatives. Although the proposed approach was inspired by Voxblox++, our results surpass its accuracy, on average, by a significant 16.1 mAP points, underscoring the impact of our improvements. Furthermore, we draw attention to the inherent pitfalls of the conventional practice of relying on ground truth camera poses as input as it introduces considerable disparities between reported outcomes and real-world performance.